# Linguistic Generalizations are not Rules: Impacts on Evaluation of LMs

### v1

@@ 8 Conclusion
Generalization is a key component of human language—and a big part of why LMs are successful at processing language. But we have argued that evaluations of the linguistic abilities of LMs are too often based on an assumption that generalization requires algebraic rules operating on words. But natural languages are not Lego sets. We suggest instead that language involves flexible combinations of rich and varied constructions of varying sizes, complexity, and degrees of abstraction, which differ from algebraic rules in many ways. By designing new evaluations that accurately reflect the complexities of language, we can avoid under- or overestimating language models. The extent to which LMs produce and interpret combinations of constructions has been limited to date. LMs offer fertile ground for new types of evaluations and new analyses that offer deeper understanding of the remarkable skills required for natural language.
@@
The claims here are based on existing evaluations of LMs. LMs are rapidly improving in a variety of ways. While we have aimed to discuss benchmarks and evaluations in ways that reflect the historical trajectory as well as the present-day landscape, it is always possible that newer models could behave differently than our characterizations.
Another potential limitation is that evaluating LMs is a moving target. While we give a number of suggestions for evaluation, we also recognize that the kinds of linguistic evaluations needed may move away from tests of grammaticality (which seems largely mastered) and towards more general kinds of language understanding. The rapid pace of LM technology makes future-proofing such designs potentially difficult.
Finally, while we focus mostly on constructionist approaches, there are related usage-based approaches from the functionalist tradition that would likely make similar predictions. Future work can further flesh out these directions.

### v2

@@ 8 Conclusion
Generalization is a key component of human language—and a big part of why LMs are successful at processing language. But we have argued that evaluations of the linguistic abilities of LMs are too often based on an assumption that generalization requires algebraic rules operating on words. Natural languages are not Lego sets. Instead, language involves flexible combinations of rich and varied constructions of differing sizes, complexities, and degrees of abstraction, which differ from algebraic rules in many ways. By designing new evaluations that accurately reflect the complexities of language, we can avoid under- or overestimating language models. The extent to which LMs produce and interpret combinations of constructions has only begun to be explored. We believe future progress lies not in asking whether LMs obey abstract rules, but in probing what kinds of constructions they learn, how they relate them, and how those structures guide novel interpretation and production. In doing so, we may better capture what it truly means to comprehend and use language.
@@
While we have aimed to discuss benchmarks and evaluations in ways that reflect the historical trajectory as well as the present-day landscape, evaluations of LMs are continually developing. We feel the dominant paradigms have and continue to be based on data generated by rules and evaluated without regard for context effects, gradience, or semantic nuance, but we are keenly aware that we have likely overlooked metrics that go beyond rule-based evaluations (e.g., Parrish et al., 2021).
We recognize the growing work in multilingual evaluations, which are inherently valuable Mueller et al. (2020); Jumelet et al. (2025); Kryvosheieva and Levy (2025). The current perspective applies to all natural languages, but comparative work is not the focus of the current perspective, and we use English examples for the sake of easy comprehension and brevity.
We have taken pains to explain the appeal of constructions over rules (Table 1 and §5), but since constructions may be less familiar, we only hope that our brief perspective may whet readers’ appetite to learn more.